# Online Housing Market

### v1

@@ Abstract
This paper studies an online variant of the celebrated housing market problem(?), where each agent has a single house and seeks to exchange it for another based on her preferences. In this online setting, agents may arrive and depart at any time, meaning that not all agents are present on the housing market simultaneously. I extend the well-known serial dictatorship and Gale’s top trading cycle mechanisms to this online scenario, aiming to retain their desirable properties such as Pareto efficiency, individual rationality, and strategy-proofness. These extensions also seek to prevent agents from strategically delaying their arrival or advancing their departure. I demonstrate that achieving all of these properties simultaneously is impossible in the online context, and I present several variants that achieve different subsets of these properties.
@@ 1 Introduction
Allocating indivisible resources to agents is a fundamental problem in computational social choice, which lies at the intersection of economics(?) and computer science(?, ?). The decision maker in such problems must consider both the preferences of the agents over the resources and their strategic behavior. In this paper, I focus on a specific problem known as the housing market in the matching theory literature(?), where each agent is endowed with a single resource that they are willing to exchange for another. Furthermore, I assume that no monetary compensation is allowed to offset any unfavorable exchanges during the process. Despite its simplicity, this problem has many applications, including the exchange of dormitory rooms for students(?), the trade of used items(?), kidney exchanges where incompatible donor/recipient pairs exchange viable organs for transplants(?), and many more (see, e.g., (?) for additional applications).
In this paper, I assume that the preferences over resources are ordinal, which is a standard assumption for the housing market. By doing so, I implicitly assume that the decision maker cannot assess the magnitude of preferences between resources, either because this information is unavailable or because it is too costly to obtain through interaction with the agents. The procedure for reallocating resources among agents is centralized, and the decision maker aims to produce an allocation that is as efficient as possible
@@
The standard offline setting assumes that all agents participating in the exchange are available at the same time, and that the exchange procedure is conducted during this period. However, in many contexts, this requirement is unrealistic or too demanding, making it difficult to gather a larger set of agents for the exchange. In more realistic scenarios, agents are only available during restricted time periods, meaning that some may not be able to participate in the exchange simultaneously. This is the case, for example, with online exchange websites(?), where agents do not arrive at the marketplace at the same time and cannot remain indefinitely before their swaps take place. This type of scenario is referred to as online(?), or dynamic, in the context of matching(?). In this paper, I show how the standard procedures, serial dictatorship and TTC, can be adapted to fit an online environment.
@@
Section 2 provides a non exhaustive list of related works, that are connected either to the housing market problem or the online version of social choice problems. Section 3 provides the main definitions of the main properties that I seek to achieve for my mechanisms. Sections 4 to 6 provide online mechanisms based on the famous serial dictatorship procedure. More specifically, Section 4 considers first a static variant of this procedure, where assignment of items to non leaving agents are irrevocable. Section 5 relax this requirement by considering a dynamic version of serial dictatorship procedure. Section 6 consider a safe version of the serial dictatorship where no agent can be allocated an item that she least prefer than her own. Finally, Section 7 introduce multiple online variants of the TTC procedure.
@@ 2.2 Online Problems in Social Choice
Multiple social problems related to my exchange problem have been examined through the lens of online procedures. In fair division, which aims to allocate resources fairly, several online variants have been explored(?, ?, ?). For instance, strategy-proofness and Pareto efficiency have been studied in this context(?), alongside envy-freeness. Various extensions of the serial dictatorship procedure have been proposed to achieve these properties(?). The online electric vehicle charging problem(?), which focuses on scheduling charging for customers arriving in an online fashion, also shares similarities with my problem. However, in both cases, the online setting differs from mine, as resources do not arrive dynamically and are known to the procedure from the outset.
The online setting has also been explored in the context of stable matching, where preferences are two-sided and pairs of agents are matched to achieve stability, with applications such as student-university allocations. The case where students or universities arrive in an online fashion has been studied(?), and mechanisms based on the standard deferred acceptance algorithm have been proposed to address this issue. Online updates to the instance may also arise from the fact that the matching process consists of multiple rounds, during which agents may unexpectedly or strategically alter their revealed preferences between rounds(?, ?, ?). In the same vein, but closer to my setting as it addresses the one-sided problem of assigning resources to agents, an online version where preferences are elicited incrementally by querying agents has been considered (?). A procedure that elicits preferences to achieve Pareto-optimality while minimizing the number of queries has been proposed.
@@
It is worth noting that there is a substantial body of literature on online algorithms (?), which aim to maximize the sum of the weights of matched pairs in an online assignment problem where weights are assigned to pairs to be matched(?). However, in my context, these approaches assume that preferences are cardinal and aim to maximize the agents’ social welfare, i.e., the sum of their utilities. Their goal is to design online algorithms with a high competitive ratio, where the competitive ratio is the smallest ratio between the value of the computed allocation and the value of the optimal offline allocation.
@@ 3 Preliminary
Let $N=\{1,\ldots,n\}$ denote the set of agents. Each agent $i$ owns a single good [^1] The resources in this paper are interchangeably referred to as goods or items $e_{i}$, called her initial endowment. Let $\mathcal{T}=[t^{-},t^{+}]$ denote the timeline during which the market is open, where $t^{-}$ and $t^{+}$ represent the opening and closing times of the market, respectively. For each agent $i$, $a_{i}$ and $d_{i}$ represent her arrival and departure times in the market, with $a_{i},d_{i}\in\mathcal{T}$ such that $t^{-}\leq a_{i}<d_{i}\leq t^{+}$. To simplify the setting, I assume that no two arrival or departure times occur simultaneously. Let $E=\{e_{1},\ldots,e_{n}\}$ represent the set of items belonging to the agents in $N$. Each agent $i$ has a strict ordinal preference $\succ_{i}$ over the items in $E$, such that $e_{j}\succ_{i}e_{k}$ means agent $i$ strictly prefers $e_{j}$ over $e_{k}$. Furthermore, symbol $\succeq_{i}$ stands for $\succ_{i}$ or $=$. An instance of the online exchange problem is represented by the set of tuples $\{(i,e_{i},a_{i},d_{i},\succ_{i})\}_{i\in N}$. Without loss of generality, I assume that the agent indices are ordered by increasing arrival times, i.e., such that $a_{1}<a_{2}<\ldots<a_{n}$ hold. Let $\mathcal{I}$ denote the entire set of possible instances.
My goal is to allocate to each agent a single item such that each item is assigned only once. In other words, I search for an allocation $M:N\rightarrow E$ such that $M(i)\neq M(j)$ for each $i\neq j$. The timeline constraints are such that each agent will leave the market with an item which belongs to an agent that arrived in the market earlier than her departure time. For any $t\in\mathcal{T}$, let $N_{<t}$ denote the subset of agents arriving before $t$, i.e., containing any agent $i\in N$ such that $a_{i}<t$. For a given instance $I\in\mathcal{I}$, an allocation $M$ is $I$-compatible if $M(i)\in E_{<d_{i}}$ holds for each agent $i$, where $E_{<t}=\{e_{i}:i\in N_{<t}\}$ for each $t\in\mathcal{T}$. For any instance $I$, the set of $I$-compatible allocations is denoted $\mathcal{M}(I)$.
An exchange algorithm $\mathcal{A}$ returns an $I$-compatible allocation for each instance $I$. By abuse of notation, I assume that $\mathcal{A}_{i}(I)$ denotes the item allocated to agent $i$ by algorithm $\mathcal{A}$ for instance $I$. In the online version of the problem, the algorithm should make a decision on the allocation for an agent when she leaves the market, without knowing the agents that arrive after her departure. In other words, the mechanism should make a decision only based on the agents that already visited the market by the departure time of the agent. For any $t\in\mathcal{T}$, let $I_{<t}$ denote a truncated copy of $I$ restricted to the agents of $N_{<t}$.
@@
The choice between online exchange algorithms should be guided by the properties fulfilled by the allocation that they compute. To compare the online exchange algorithms, I consider multiple standard desiderata properties that could be considered as desirable or necessary. The first one is a standard definition of efficiency in multi-agent decision problems.
@@
The individual rationality property belongs to the set of properties that incentivize agents to participate in the exchange algorithm, as no agent will receive an item that is less desirable than her initial endowment. Another standard property related to the manipulative power of agents, through misreporting their preferences, is called incentive compatibility. I generalize the notion of incentive compatibility to also take into account manipulations related to arrival and departure times.
@@
Note that the notion denoted $WIC$ in my setting corresponds to the standard notion of incentive compatibility in the offline setting. Note also that I implicitly assumed that the manipulative power of an agent concerning her arrival and departure time is restricted, since she cannot claim an arrival time earlier than her true arrival time, nor a departure time later than her true departure time. This assumption is made because I do not believe that this type of misreporting is realistic in an exchange market. Furthermore, my goal is to push agents to stay as long as possible in the market to obtain a better allocation.
@@ 4 Static Serial Dictatorship Procedure
In this section, I consider an exchange algorithm based on the serial dictatorship procedure, which is described in Algorithm 1. The standard version of the serial dictatorship procedure is based on a permutation of the agents defining the order in which each agent will choose her item among the remaining ones. I slightly generalize these permutations by considering permutation functions whose order depends on the instance. More formally, $\Pi:\mathcal{I}\rightarrow N^{N}$ denotes this permutation function, which is such that for any instance $I$ and for any $i\in\{1,\ldots,n\}$, $\Pi_{i}(I)$ denotes the agent choosing at position $i$. Despite the fact that this permutation theoretically may depend on the full instance, I assume in this paper that it only depends on the arrival and departure times of the agents. An example of such a permutation is the ascending departure permutation $\delta$ which ranks the agents according to their departure times, [^3] The input instance is omitted to simplify notations. and more specifically by increasing departure times, i.e., such that $d_{\delta(1)}<d_{\delta(2)}<\ldots<d_{\delta(n)}$.
@@
In Algorithm 1, every iteration of the ”for” loop corresponds to the departure of an agent, specifically agent $\delta(i)$, whose departure time is the $i^{th}$ earliest. However, the order in which the agents choose their items is determined by the permutation function $\Pi$. Therefore, before agent $\delta(i)$ chooses her item, all agents that are ranked before her according to $\Pi$ must choose their items. In this static (or greedy) version of the algorithm, once an agent has chosen an item, she is permanently matched to it, even if more desirable items arrive later. Note that the procedure $best_{i}$ returns the most preferred item of agent $i$ from a given set of items. Algorithm 1 is designed to be online, as only the part of the instance corresponding to agents who have already arrived, $I_{<d_{\delta(i)}}$, is used at iteration $i$, during the departure of agent $\delta(i)$. However, the permutation used as input must be consistent with this online algorithm in the sense that the order of agents already matched to items must not change when new agents arrive. Otherwise, an agent might be matched multiple times or ignored by the algorithm.
@@
The permutation function $\Pi$ is static online consistent ($SOC$), if for each agent $i$, and for any positions $i^{\prime}$ and $j^{\prime}$ such that $\Pi_{i^{\prime}}(I)=i$ and $j^{\prime}\leq i^{\prime}$, $\Pi_{j^{\prime}}(I)=\Pi_{j^{\prime}}(I_{<d_{i}})$ holds.
The following proposition shows that Algorithm 1 is online and weakly incentive compatible when the permutation used is $SOC$.
@@
Algorithm 1 is both online and $WIC$ if the input permutation function is $SOC$.
@@
Let me first show that it is online. Let $I$ be an instance and $i$ be an agent. I need to show that Algorithm 1 applied to both $I$ and $I_{<d_{i}}$ returns the same outcome to agent $i$. To simplify notations, I denote by $I^{\prime}$ instance $I_{<d_{i}}$. Let $i^{\prime}$ denote the position of agent $i$ in $\Pi(I)$, i.e., $\Pi_{i^{\prime}}(I)=i$ holds. Note first that since $\Pi$ is $SOC$, I know that for any position $j^{\prime}$ such that $j^{\prime}\leq i^{\prime}$, I have $\Pi_{j^{\prime}}(I)=\Pi_{j^{\prime}}(I^{\prime})$. For the same reason, I know that for any agent $j$ and positions $j^{\prime}$ and $k^{\prime}$ such that $k^{\prime}\leq j^{\prime}\leq i^{\prime}$ and $\Pi_{j^{\prime}}(I)=\Pi_{j^{\prime}}(I^{\prime})=j$, I have $\Pi_{k^{\prime}}(I)=\Pi_{k^{\prime}}(I_{<d_{j}})$ and $\Pi_{k^{\prime}}(I^{\prime})=\Pi_{k^{\prime}}(I^{\prime}_{<d_{j}})$. But since $I^{\prime}=I_{<d_{i}}$, this implies $I_{<d_{j}}=I^{\prime}_{<d_{j}}$ and $\Pi_{k^{\prime}}(I_{<d_{j}})=\Pi_{k^{\prime}}(I^{\prime}_{<d_{j}})$ hold for any $k^{\prime}\leq j^{\prime}\leq i^{\prime}$. Therefore, the first iterations of the “for” loop of Algorithm 1, until agent $i$ leaves the market at time $d_{i}$, makes the same assignment to the same agents for $I$ and $I^{\prime}$. Therefore, agent $i$ receives the exact same item in both cases.
I now show that it is $WIC$. The proof relies on the fact that the standard serial dictatorship procedure is incentive compatible in the offline setting. This is due to the fact that the preferences of an agent are only considered once to assign to this agent her most favorite item among the remaining ones (by function $best$). Therefore, an agent has no incentive to misreport her preferences since the item that she will receive cannot be more preferred by doing so. ∎

### v2

@@ Abstract
This paper studies an online variant of the celebrated housing market problem (?), where each agent owns a single house and seeks to exchange it for another based on her preferences. In this online setting, agents may arrive and depart at any time, meaning that not all agents are present on the housing market simultaneously. We extend the well-known serial dictatorship and Gale’s top trading cycle mechanisms to this online scenario, aiming to retain their desirable properties such as Pareto efficiency, individual rationality, and strategy-proofness. These extensions also seek to prevent agents from strategically delaying their arrival or advancing their departure. We demonstrate that achieving all of these properties simultaneously is impossible in the online context, and we present several variants that achieve different subsets of these properties.
@@ 1 Introduction
Allocating indivisible resources to agents is a fundamental problem in computational social choice, which lies at the intersection of economics (?) and computer science (?, ?). The decision maker in such problems must consider both the preferences of the agents over the resources and their strategic behavior. In this paper, we focus on a specific problem known as the housing market in the matching theory literature (?), where each agent is endowed with a single resource that they are willing to exchange for another. Furthermore, we assume that no monetary compensation is allowed to offset any unfavorable exchanges during the process. Despite its simplicity, this problem has many applications, including the exchange of dormitory rooms for students (?), the trade of used items (?), kidney exchanges where incompatible donor/recipient pairs exchange viable organs for transplants (?), and many more (see, e.g., (?) for additional applications).
In this paper, we assume that the preferences over resources are ordinal, which is a standard assumption for the housing market. By doing so, we implicitly assume that the decision maker cannot assess the magnitude of preferences between resources, either because this information is unavailable or because it is too costly to obtain through interaction with the agents. The procedure for reallocating resources among agents is centralized, and the decision maker aims to produce an allocation that is as efficient as possible
@@
The standard offline setting assumes that all agents participating in the exchange are available at the same time, and that the exchange procedure is conducted during this period. However, in many contexts, this requirement is unrealistic or too demanding, making it difficult to gather a larger set of agents for the exchange. In more realistic scenarios, agents are only available during restricted time periods, meaning that some may not be able to participate in the exchange simultaneously. This is the case, for example, with online exchange websites (?), where agents do not arrive at the marketplace at the same time and cannot remain indefinitely before their swaps take place. This type of scenario is referred to as online (?), or dynamic, in the context of matching (?). In this paper, we show how the standard procedures, serial dictatorship and TTC, can be adapted to fit an online environment.
@@
Section 2 provides a non-exhaustive list of related works connected either to the housing market problem or to online versions of social choice problems. Section 3 presents the main definitions of the key properties that we seek to achieve for our mechanisms. Sections 4 to 6 describe online mechanisms based on the famous serial dictatorship procedure. More specifically, Section 4 first considers a static variant of this procedure, where assignments to non-leaving agents are irrevocable. Section 5 relaxes this requirement by considering a dynamic version of the serial dictatorship. Section 6 considers a safe version in which no agent can be allocated an item she prefers less than her own. Finally, Section 7 introduces multiple online variants of the TTC procedure.
@@ 2.2 Online Problems in Social Choice
Multiple social problems related to our exchange problem have been examined through the lens of online procedures. In fair division, which aims to allocate resources fairly, several online variants have been explored (?, ?, ?). For instance, strategy-proofness and Pareto efficiency have been studied in this context (?), alongside envy-freeness. Various extensions of the serial dictatorship procedure have been proposed to achieve these properties (?). The online electric vehicle charging problem (?), which focuses on scheduling charging for customers arriving in an online fashion, also shares similarities with our problem. However, in both cases, the online setting differs from mine, as resources do not arrive dynamically and are known to the procedure from the outset.
The online setting has also been explored in the context of stable matching, where preferences are two-sided and pairs of agents are matched to achieve stability, with applications such as student-university allocations. The case where students or universities arrive in an online fashion has been studied (?), and mechanisms based on the standard deferred acceptance algorithm have been proposed to address this issue. Online updates to the instance may also arise from the fact that the matching process consists of multiple rounds, during which agents may unexpectedly or strategically alter their revealed preferences between rounds (?, ?, ?). In the same vein, but closer to our setting as it addresses the one-sided problem of assigning resources to agents, an online version where preferences are elicited incrementally by querying agents has been considered (?). A procedure that elicits preferences to achieve Pareto-optimality while minimizing the number of queries has been proposed.
@@
It is worth noting that there is a substantial body of literature on online algorithms (?), which aim to maximize the sum of the weights of matched pairs in an online assignment problem where weights are assigned to pairs to be matched (?). However, in our context, these approaches assume that preferences are cardinal and aim to maximize the agents’ social welfare, i.e., the sum of their utilities. Their goal is to design online algorithms with a high competitive ratio, where the competitive ratio is the smallest ratio between the value of the computed allocation and the value of the optimal offline allocation.
@@ 3 Preliminary
Let $N=\{1,\ldots,n\}$ denote the set of agents. Each agent $i$ owns a single good [^1] The resources in this paper are interchangeably referred to as goods or items $e_{i}$, called her initial endowment. Let $\mathcal{T}=[t^{-},t^{+}]$ denote the timeline during which the market is open, where $t^{-}$ and $t^{+}$ represent the opening and closing times of the market, respectively. For each agent $i$, $a_{i}$ and $d_{i}$ represent her arrival and departure times in the market, with $a_{i},d_{i}\in\mathcal{T}$ such that $t^{-}\leq a_{i}<d_{i}\leq t^{+}$. To simplify the setting, we assume that no two arrival or departure times occur simultaneously. [^2] Note that when the exchange market takes place in a web-based application, times are recorded precisely, and equal timestamps are highly unlikely. Furthermore, if such an event occurs, random tie-breaking rules may be used. Let $E=\{e_{1},\ldots,e_{n}\}$ represent the set of items belonging to the agents in $N$. Each agent $i$ has a strict ordinal preference $\succ_{i}$ over the items in $E$, such that $e_{j}\succ_{i}e_{k}$ means agent $i$ strictly prefers $e_{j}$ over $e_{k}$. Furthermore, symbol $\succeq_{i}$ stands for $\succ_{i}$ or $=$. An instance of the online exchange problem is represented by the set of tuples $\{(i,e_{i},a_{i},d_{i},\succ_{i})\}_{i\in N}$. Without loss of generality, we assume that the agent indices are ordered by increasing arrival times, i.e., such that $a_{1}<a_{2}<\ldots<a_{n}$ hold. Let $\mathcal{I}$ denote the entire set of possible instances.
Our goal is to allocate to each agent a single item such that each item is assigned only once. In other words, we search for an allocation $M:N\rightarrow E$ such that $M(i)\neq M(j)$ for each $i\neq j$. The timeline constraints are such that each agent will leave the market with an item which belongs to an agent that arrived in the market earlier than her departure time. For any $t\in\mathcal{T}$, let $N_{<t}$ denote the subset of agents arriving before $t$, i.e., containing any agent $i\in N$ such that $a_{i}<t$. For a given instance $I\in\mathcal{I}$, an allocation $M$ is $I$-compatible if $M(i)\in E_{<d_{i}}$ holds for each agent $i$, where $E_{<t}=\{e_{i}:i\in N_{<t}\}$ for each $t\in\mathcal{T}$. For any instance $I$, the set of $I$-compatible allocations is denoted $\mathcal{M}(I)$.
An exchange algorithm $\mathcal{A}$ returns an $I$-compatible allocation for each instance $I$. By abuse of notation, we assume that $\mathcal{A}_{i}(I)$ denotes the item allocated to agent $i$ by algorithm $\mathcal{A}$ for instance $I$. In the online version of the problem, the algorithm should make a decision on the allocation for an agent when she leaves the market, without knowing the agents that arrive after her departure. In other words, the mechanism should make a decision only based on the agents that already visited the market by the departure time of the agent. For any $t\in\mathcal{T}$, let $I_{<t}$ denote a truncated copy of $I$ restricted to the agents of $N_{<t}$.
@@
The choice between online exchange algorithms should be guided by the properties fulfilled by the allocation that they compute. To compare the online exchange algorithms, we consider multiple standard desiderata properties that could be considered as desirable or necessary. The first one is a standard definition of efficiency in multi-agent decision problems.
@@
The individual rationality property belongs to the set of properties that incentivize agents to participate in the exchange algorithm, as no agent will receive an item that is less desirable than her initial endowment. Another standard property related to the manipulative power of agents, through misreporting their preferences, is called incentive compatibility. We generalize the notion of incentive compatibility to also take into account manipulations related to arrival and departure times.
@@
In Definition 4, instance $I$ refers to the setting in which all agents truthfully report their preferences and arrival/departure times, while $I^{\prime}$ is identical except that agent $i$ may misreport her preferences and/or availability. We aim for mechanisms in which such misreports do not yield a more favorable outcome for agent $i$. The notion denoted $WIC$ in our setting, where agent $i$ cannot misreport her arrival or departure time, corresponds to the standard notion of incentive compatibility in the offline setting. The arrival time $a_{i}$ represents the earliest possible time at which agent $i$ can enter the market—she cannot arrive earlier due to external constraints (e.g., unavailability, continued use of the item). Similarly, $d_{i}$ denotes the latest possible departure time. This is why agent $i$ cannot arrive before $a_{i}$ or stay beyond $d_{i}$. [^4] We implicitly assume it is possible to verify an agent’s presence in the market during her declared interval. For instance, arrival can be confirmed by collecting the item she offers, and continued presence until $d_{i}$ can be enforced by delivering her assigned item only at that time. However, agent $i$ may still strategically choose to arrive later than $a_{i}$ or leave earlier than $d_{i}$ in an attempt to manipulate the outcome. Our goal is to design mechanisms that incentivize agents to remain in the market throughout their true availability window, from $a_{i}$ to $d_{i}$. The different notions of incentive compatibility aim to ensure that agents are incentivized to arrive as early as possible ($a$-IC), as late as possible ($d$-IC), or both (SIC).
@@ 4 Static Serial Dictatorship Procedure
In this section, we consider an exchange algorithm based on the serial dictatorship procedure, which is described in Algorithm 1. The standard version of the serial dictatorship procedure is based on a permutation of the agents defining the order in which each agent will choose her item among the remaining ones. We slightly generalize these permutations by considering permutation functions whose order depends on the instance. More formally, $\Pi:\mathcal{I}\rightarrow N^{N}$ denotes this permutation function, which is such that for any instance $I$ and for any $i\in\{1,\ldots,n\}$, $\Pi_{i}(I)$ denotes the agent choosing at position $i$. Despite the fact that this permutation theoretically may depend on the full instance, we assume in this paper that it only depends on the arrival and departure times of the agents. An example of such a permutation is the ascending departure permutation $\delta$ which ranks the agents according to their departure times, and more specifically by increasing departure times, i.e., such that $d_{\delta(1)}<d_{\delta(2)}<\ldots<d_{\delta(n)}$. [^5] Note that, to simplify notation, we omit the input instance in the notation of the concrete permutation functions ($\delta$ and $\alpha$, as well as the permutation $\pi$ appearing in Algorithms 1, 2, and 3). Moreover, an agent’s position in these permutations is indicated in parentheses rather than as a subscript.
@@
In Algorithm 1, every iteration of the ”for” loop corresponds to the departure of an agent, specifically agent $\delta(i)$, whose departure time is the $i^{th}$ earliest. However, the order in which the agents choose their items is determined by the permutation function $\Pi$. Therefore, before agent $\delta(i)$ chooses her item, all agents that are ranked before her according to $\Pi$ must choose their items. In this static (or greedy) version of the algorithm, once an agent has chosen an item, she is permanently matched to it, even if more desirable items arrive later. Note that the procedure $best_{i}$ returns the most preferred item of agent $i$ from a given set of items. Algorithm 1 is designed to be online, as only the part of the instance corresponding to agents who have already arrived, $I_{<d_{\delta(i)}}$, is used at iteration $i$, during the departure of agent $\delta(i)$. However, the permutation must be prefix-consistent i.e., the order of agents already matched to items must not change when new agents arrive. Otherwise, an agent might be matched multiple times or ignored by the algorithm.
@@
Permutation function $\Pi$ is prefix-stable ($PS$), if for each agent $i$, and for any positions $i^{\prime}$ and $j^{\prime}$ such that $\Pi_{i^{\prime}}(I)=i$ and $j^{\prime}\leq i^{\prime}$, $\Pi_{j^{\prime}}(I)=\Pi_{j^{\prime}}(I_{<d_{i}})$ holds.
This definition stipulates that, for a permutation function to be prefix-stable, any agents ranked before agent $i$ when she leaves the market should retain their positions afterward. The following proposition shows that Algorithm 1 is online and weakly incentive compatible when the permutation used is prefix-stable.
@@
Algorithm 1 is both online and $WIC$ if the input permutation function is $PS$.
@@
Let us first show that it is online. Let $I$ be an instance and $i$ be an agent. We need to show that Algorithm 1 applied to both $I$ and $I_{<d_{i}}$ returns the same outcome to agent $i$. To simplify notations, we denote by $I^{\prime}$ instance $I_{<d_{i}}$. Let $i^{\prime}$ denote the position of agent $i$ in $\Pi(I)$, i.e., $\Pi_{i^{\prime}}(I)=i$ holds. Note first that since $\Pi$ is $PS$, we know that for any position $j^{\prime}$ such that $j^{\prime}\leq i^{\prime}$, we have $\Pi_{j^{\prime}}(I)=\Pi_{j^{\prime}}(I^{\prime})$. For the same reason, we know that for any agent $j$ and positions $j^{\prime}$ and $k^{\prime}$ such that $k^{\prime}\leq j^{\prime}\leq i^{\prime}$ and $\Pi_{j^{\prime}}(I)=\Pi_{j^{\prime}}(I^{\prime})=j$, we have $\Pi_{k^{\prime}}(I)=\Pi_{k^{\prime}}(I_{<d_{j}})$ and $\Pi_{k^{\prime}}(I^{\prime})=\Pi_{k^{\prime}}(I^{\prime}_{<d_{j}})$. But since $I^{\prime}=I_{<d_{i}}$, this implies $I_{<d_{j}}=I^{\prime}_{<d_{j}}$ and $\Pi_{k^{\prime}}(I_{<d_{j}})=\Pi_{k^{\prime}}(I^{\prime}_{<d_{j}})$ hold for any $k^{\prime}\leq j^{\prime}\leq i^{\prime}$. Therefore, the first iterations of the “for” loop of Algorithm 1, until agent $i$ leaves the market at time $d_{i}$, makes the same assignment to the same agents for $I$ and $I^{\prime}$. Therefore, agent $i$ receives the exact same item in both cases.
We now show that it is $WIC$. The proof relies on the fact that the standard serial dictatorship procedure is incentive compatible in the offline setting. This is due to the fact that the preferences of an agent are only considered once to assign to this agent her most favorite item among the remaining ones (by function $best$). Therefore, an agent has no incentive to misreport her preferences since the item that she will receive cannot be more preferred by doing so. ∎